# Bethe equations for the critical three-state Potts spin chain with toroidal boundary conditions

### v1

@@ 1 Introduction
$\left[T_{\mathrm{dia}}(x)\right]^{b_{1},\dots,b_{L}}_{a_{1},\dots,a_{L}}=\prod_{j=1}^{L}W_{v}(a_{j},b_{j}|x)W_{h}(a_{j},b_{j+1}|x),$ (3)
@@
We have organized this paper as follows. In the next section, we introduce toroidal boundary conditions for an integrable $n$-state spin model by exploring a recent mapping to an equivalent solvable $n$-state vertex model [19]. This permits us to formulate twisted boundary conditions in terms of boundary seams, responsible for changing the interactions at one end of the respective quantum spin chains. We apply this construction to the three-state scalar Potts model, making it possible to build the transfer matrices whose Hamiltonian limits give rise to the quantum spin chains $H^{(\pm)}$ and $H^{(c)}$ discussed above. In sections 3 and 4, we used certain properties of the transfer matrices eigenvalues to establish the Bethe equations parameterizing the spectra of the Hamiltonians $H^{(+)}$ and $H^{(c)}$. We investigate the solutions of the Bethe equations for $L=2$ and we use the Bethe roots to compute the Hamiltonian energies and the momenta of the eigenstates. In section 5, we use our framework to build integrable Hamiltonians with periodic boundary conditions whose spectra are obtained by combining the eigenvalues of the spin chains with different types of toroidal boundary conditions. In section 6, we present our concluding remarks. In the appendices $A$ and $B$, we summarize the analysis concerning the completeness of the solutions of the Bethe equations for $L=3$.
@@ 2 Toroidal Boundary Conditions
We start this section by describing the main results concerning an equivalence among integrable $n$-state spin and vertex models on the square lattice [19]. We recall that such a mapping assumes that the horizontal $W_{h}(i,j|x)$ and vertical $W_{v}(i,j|x)$ edge weights of the spin model are parametrized in terms of the difference of spectral parameters. It is well known that an important ingredient for the formulation of the integrability of a given vertex model is the corresponding Lax operator [20]. In this sense, it has been argued that such an operator can be constructed from the spin model edge weights as follows [19],
$\mathbb{L}_{12}(x)=\sum_{i,j,k=1}^{n}W_{h}(j,i|x)W_{v}(j,k|x)e_{ik}\otimes e_{ji},$ (17)
where $e_{i,j}$ denotes the standard $n\times n$ Weyl matrices.
@@
$\mathrm{T}(x)=\mathrm{Tr}_{\cal{A}}\left[\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}L}(x)\mathbb{L}_{{\cal{A}}L-1}(x)\dots\mathbb{L}_{{\cal{A}}1}(x)\right],$ (21)
where the symbol $\cal{A}$ denotes the auxiliary space $\mathbb{C}^{n}$. We note that the periodic boundary condition corresponds to choosing $\mathrm{G}$ as the $n\times n$ identity matrix.
@@
At this point, we also note that $\mathrm{T}(0)$ plays the role of a translation operator which shifts the spin generators by one site modulo $L$. More precisely, one can verify the following relations,
$\mathrm{T}(0)\left[\mathrm{H}^{(\mathrm{bulk})}_{j,j+1}\right]\mathrm{T}^{-1}(0)=\mathrm{H}^{(\mathrm{bulk})}_{j+1,j+2},~~j=1,\dots,L-2,$
$\mathrm{T}(0)\left[\mathrm{H}^{(\mathrm{bulk})}_{L-1,L}\right]\mathrm{T}^{-1}(0)=\mathrm{H}^{(\mathrm{bound})}_{L,1}$. (25)
@@
$\mathrm{T}^{(\pm)}(x)=\mathrm{Tr}_{{\cal{A}}}\left[\mathrm{G}^{(\pm)}_{{\cal A}}\mathbb{L}_{{\cal A}L}(x)\mathbb{L}_{{\cal A}L-1}(x)\dots\mathbb{L}_{{\cal A}1}(x)\right],~~~\mathrm{T}^{(c)}(x)=\mathrm{Tr}_{{\cal{A}}}\left[\mathrm{G}^{(c)}_{\cal{A}}\mathbb{L}_{{\cal A}L}(x)\mathbb{L}_{{\cal A}L-1}(x)\dots\mathbb{L}_{{\cal A}1}(x)\right].$ (27)
The transfer matrices $\mathrm{T}^{(\pm)}(x)$ are invariant under the $Z(3)$ spin rotation and their Hamiltonian limits, apart from a constant bulk term, correspond to the quantum spin chains $H^{(\pm)}$. Thus, we have the following commutation relations,
$\mathrm{T}^{(\pm)}(x),{\cal{O}}_{Z(3)}]=0,~~[\mathrm{T}^{(\pm)}(x),H^{(\pm)}]=0$ (28)
such that the spectra of $\mathrm{T}^{(+)}(x)$ and $\mathrm{T}^{(-)}(x)$ are related by the following equivalences,
$\mathrm{Spec}\big[T^{(+)}\big]_{Q=0}=\mathrm{Spec}\big[T^{(-)}\big]_{Q=0},~~\mathrm{Spec}\big[T^{(+)}\big]_{Q=1}=\mathrm{Spec}\big[T^{(-)}\big]_{Q=2},~~\mathrm{Spec}\big[T^{(+)}\big]_{Q=2}=\mathrm{Spec}\big[T^{(-)}\big]_{Q=1}$ (29)
The transfer matrix $\mathrm{T}^{(c)}(x)$ keeps only the invariant concerning the $Z(2)$ subgroup associated with the charge conjugation. The respective Hamiltonian limit gives rise to the quantum spin chain $H^{(c)}$, namely
$\mathrm{T}^{(c)}(x),{\cal{O}}_{Z(2)}]=0,~~[\mathrm{T}^{(c)}(x),H^{(c)}]=0$ (30)
In the next sections, we shall see that the construction of these transfer matrices is essential to set up the Bethe equations for the corresponding quantum spin chains. From the spectral relations (29) it is enough to consider the problem for $T^{(+)}(x)$ and $T^{(c)}(x)$.
3 The Bethe equation for $T^{(+)}(x)$
Here, we follow the strategy devised by Baxter to solve the eigenvalue problem associated with the diagonal-to-diagonal transfer matrix of the Ising model [1]. We recall that this kind of approach has also been used to uncover the Bethe equation for the critical $Z(n)$ spin model with periodic boundary conditions [11, 22]. The first step is to build a suitable proposal for the form of the eigenvalues of the transfer matrix $T^{(+)}$. To this end, we observe any matrix elements of such operator have the following structure,
$\left[T^{(+)}\right]_{a_{1},\dots,a_{L}}^{b_{1},\dots,b_{L}}=\Big[\frac{1}{g(x)g_{1}(x)}\Big]^{L}H(x)$ (31)
@@ 3 The Bethe equation for $T^{(+)}(x)$
$T^{(+)}(x)\left|\phi\right\rangle=\Lambda^{(+)}(x)\left|\phi\right\rangle$ (34)
@@
Due to the independence of the eigenstates on $x$ a given eigenvalue $\Lambda^{(+)}(x)$ can be regarded as a linear combination of the elements of the transfer matrix $T^{(+)}(x)$ with constant coefficients. As a consequence, we can conclude that $\Lambda^{(+)}(x)$ should have the same form as given by Eqs.(31,33). To obtain further insights on the structure of $\Lambda^{(+)}(x)$ we have studied the eigenvalue problem (34) for small finite sizes. This analysis reveals to us that the the form of eigenvalues depends on the $Z(3)$ sectors besides not given solely in terms of the product of sines. Investigating the asymptotic behavior of $\Lambda^{(+)}(x)$ for $x\rightarrow\pm i\infty$ leads to propose the following form,
@@
The next step is to determine the set of nonlinear equations satisfied by the variables $\xi_{k}$ on a finite interval. This task can be done by exploring certain matrix identity among the products of transfer matrices with suitable distinct spectral parameters. Here we have been inspired by the matrix structure derived for the diagonal-to-diagonal transfer matrix of the three-state Chiral Potts model [12, 13]. We have found that the transfer matrix $T^{(+)}(x)$ satisfies the following identity,
$T^{(+)}(x-\frac{\pi}{3})T^{(+)}(x-\frac{\pi}{6})T^{(+)}(x)$ $=$ $T^{(+)}(0)\bigg\{\big[f_{1}(x)\big]^{L}T^{(+)}(x-\frac{\pi}{3})+\big[f_{2}(x)\big]^{L}T^{(+)}(x)$ (37)
$+$ $\big[f_{3}(x)\big]^{L}T^{(+)}(x+\frac{\pi}{3})\bigg\}$
@@
At this point, we note that Eqs.(40,41) leads us to the expressions of the Bethe equation (11) and the energy eigenvalue (12) presented in the introduction by choosing,
@@ 3.1 The Bethe roots for $L=2$
We now describe the set of roots solving the Bethe equation (11) associated with the energy eigenvalues of the Hamiltonian $H^{(+)}$ for $L=2$. We obtain some insight on the nature of the roots, exploring the fact that the eigenvectors of $T^{(+)}(x)$ are independent of the spectral parameter. By choosing a suitable value for $x$ we can compute numerically the eigenvectors making it possible the determination of the corresponding eigenvalues as polynomial functions on $x$. The zeros of such polynomials work as starting values for the numerical solution of the Bethe equation (11) providing us the Bethe roots with high precision.
@@
In table (1) we have summarized our results for $L=2$ considering the total number of nine eigenstates. For each energy of the spectrum of $H^{(+)}$ we exhibit the sector $Q$, the spin $s_{p}$ of the state and the corresponding roots $\{\lambda_{j}\}$ solving the Bethe equation (11). The energy values are obtained with the help of Eqs.(12,13) and we have verified that they agree with those obtained from the exact diagonalization of the Hamiltonian $H^{(+)}$. From table (1) we note that if a given energy in sector $Q=1$ is parametrized by a set of roots $\{\lambda_{j}\}$ the same energy is obtained in sector $Q=2$ using the opposite set of roots $\{-\lambda_{j}\}$. From our solution, we see that this property is valid for arbitrary $L$ in accordance with the fact that the spectra for sectors $Q=1,2$ are the same. In addition, from Eqs.(13,43) we see that the momenta of such states are however opposite to each other.
$E^{(+)}$ $Q$ $s_{p}$ $\lambda_{j}$
-4.93624921 0 0 $\pm 0.53202156i$
-2.30940107 1 $\frac{1}{3}$ $0.06315064\pm 0.54254935i,~0.19822180$
-2.30940107 2 $\frac{1}{3}$ $0.06315064\pm 0.54254935i$0.19822180-0.19822180
-0.84529946 1 $\frac{2}{3}$ $0.16532518\pm 0.50188873i$0.10277717-0.10277717
-0.84529946 2 $\frac{2}{3}$ $0.16532518\pm 0.50188873i,~0.10277717$
1.15470053 0 1 $0,~\frac{i\pi}{2}$
3.15470053 1 $\frac{2}{3}$ 0.245380170.245380170.02638725-0.026387250.64959867-0.64959867
3.15470053 2 $\frac{2}{3}$ 0.24538017-0.245380170.02638725,0.649598670.02638725,~0.64959867
3.78154867 0 0 $\pm 0.12258177$
@@
Table 1: The eigenvalues ($E^{(+)}$) of the Hamiltonian $H^{(+)}$ (10) for $L=2$. For each energy, we indicate the sector ($Q$), the spin of the state $(s_{p})$ and the values of the Bethe roots ($\lambda_{j}$).
@@
4 The Bethe equation for $T^{(c)}(x)$
Here, we use the same procedure explained in the previous section to determine the parameterization of the spectrum of Hamiltonian $H^{(c)}$ in terms of a Bethe ansatz equation. We therefore avoid the repetition of technical details and concentrate our presentation to the main results. In particular, we find out that the structure of the eigenvalues of $T^{(c)}(x)$ can be given as follows,

### v2

@@ 1 Introduction
$\left[T_{dia}(x)\right]^{b_{1},\dots,b_{L}}_{a_{1},\dots,a_{L}}=\prod_{j=1}^{L}W_{v}(a_{j},b_{j}|x)W_{h}(a_{j},b_{j+1}|x),$ (3)
@@
We have organized this paper as follows. In the next section, we introduce toroidal boundary conditions for an integrable $n$-state spin model by exploring a recent mapping to an equivalent solvable $n$-state vertex model [19]. This permits us to formulate twisted boundary conditions in terms of boundary seams, responsible for changing the interactions at one end of the respective quantum spin chains. We apply this construction to the three-state scalar Potts model, making it possible to build the transfer matrices whose Hamiltonian limits give rise to the quantum spin chains $H^{(\pm)}$ and $H^{(c)}$ discussed above. In sections 3 and 4, we used certain properties of the transfer matrices eigenvalues to establish the Bethe equations parameterizing the spectra of the Hamiltonians $H^{(+)}$ and $H^{(c)}$. We investigate the solutions of the Bethe equations for $L=2$ and we use the Bethe roots to compute the Hamiltonian energies and the momenta of the eigenstates. The respective data for the eigenvalues of the transfer matrix operators are also presented.
In section 5, we use our framework to build integrable Hamiltonians with periodic boundary conditions whose spectra are obtained by combining the eigenvalues of the spin chains with different types of toroidal boundary conditions. In section 6, we present our concluding remarks. In the appendices $A$ and $B$, we summarize the analysis concerning the completeness of the solutions of the Bethe equations for $L=3$.
@@ 2 Toroidal Boundary Conditions
We start this section by describing the main results concerning an equivalence among integrable $n$-state spin and vertex models on the square lattice [19]. We recall that such a mapping assumes that the horizontal $W_{h}(i,j|x)$ and vertical $W_{v}(i,j|x)$ edge weights of the spin model are parametrized in terms of the difference of spectral parameters. It is well known that an important ingredient for the formulation of the integrability of a given vertex model is the corresponding Lax operator [20]. This operator contains the information about the structure of the Boltzmann weights of the vertex model denoted here by $\mathbb{L}_{i,j}^{k,l}(x)$. The indices $i,j,k,l$ represent the four state variables assigned to the links of the square lattice meeting together at a given vertex. In this sense, it has been argued that such an operator can be constructed from the spin model edge weights as follows [19],
$\mathbb{L}_{12}(x)=\sum_{i,j,k,l=1}^{n}\mathbb{L}_{i,j}^{k,l}(x)~e_{ik}\otimes e_{jl}=\sum_{i,j,k,l=1}^{n}W_{h}(j,i|x)W_{v}(j,k|x)\delta_{i,l}~e_{ik}\otimes e_{jl},$ (17)
where $e_{i,j}$ denotes the standard $n\times n$ Weyl matrices and the Lax operator is schematically illustrated in Fig.(1).
[…]
Figure 1: The illustration of the elements of the Lax operator (17) on the square lattice.
@@
$T_{ver}(x)=\mathrm{Tr}_{\cal{A}}\left[\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}L}(x)\mathbb{L}_{{\cal{A}}L-1}(x)\dots\mathbb{L}_{{\cal{A}}1}(x)\right],$ (21)
where the symbol $\cal{A}$ denotes the auxiliary space $\mathbb{C}^{n}$. We note that the periodic boundary condition corresponds to choosing $\mathrm{G}$ as the $n\times n$ identity matrix. The matrix elements $T_{ver}(x)]_{a_{1}\dots a_{L}}^{b_{1}\dots b_{L}}$ of the transfer matrix are illustrated in Fig.(2) where $G(a,b)$ denotes the entries of the boundary seam $G$.
$\cdots$$T_{ver}(x)]_{a_{1}\dots a_{L}}^{b_{1}\dots b_{L}}$$\sum_{c_{1},\dots,c_{L+1}=1}^{n}$$a_{1}$$a_{2}$$a_{L-1}$$a_{L}$$b_{1}$$b_{2}$$b_{L-1}$$b_{L}$$c_{1}$$c_{2}$$c_{3}$$c_{L-1}$$c_{L}$$c_{L+1}$$G(c_{L+1},c_{1})$
Figure 2: Schematic representation of the elements of the row-to-row transfer matrix of vertex models with twisted boundary $G$.
As usual the partition function $Z_{ver}(L)$ of the vertex model on the $L\times L$ square lattice is given by the trace of the $L$-th power of the row-to-row transfer matrix,
$Z_{ver}(L)=\mathrm{Tr}_{{\cal V}}\left[\Big(T_{ver}(x)\Big)^{L}\right].$ (22)
where ${\cal V}=\displaystyle\prod_{j=1}^{L}\otimes{\mathrm{C}}^{n}$ is Hilbert space associated to the underlying quantum spin chain.
@@
At this point, we also note that $T_{ver}(0)$ plays the role of a translation operator which shifts the spin generators by one site modulo $L$. More precisely, one can verify the following relations,
$T_{ver}(0)\left[\mathrm{H}^{(\mathrm{bulk})}_{j,j+1}\right]\left[T_{ver}(0)\right]^{-1}=\mathrm{H}^{(\mathrm{bulk})}_{j+1,j+2},~~j=1,\dots,L-2,$
$T_{ver}(0)\left[\mathrm{H}^{(\mathrm{bulk})}_{L-1,L}\right]\left[T_{ver}(0)\right]^{-1}=\mathrm{H}^{(\mathrm{bound})}_{L,1}$. (26)
@@
$T_{ver}^{(\pm)}(x)=\mathrm{Tr}_{{\cal{A}}}\left[\mathrm{G}^{(\pm)}_{{\cal A}}\mathbb{L}_{{\cal A}L}(x)\mathbb{L}_{{\cal A}L-1}(x)\dots\mathbb{L}_{{\cal A}1}(x)\right],~~T_{ver}^{(c)}(x)=\mathrm{Tr}_{{\cal{A}}}\left[\mathrm{G}^{(c)}_{\cal{A}}\mathbb{L}_{{\cal A}L}(x)\mathbb{L}_{{\cal A}L-1}(x)\dots\mathbb{L}_{{\cal A}1}(x)\right].$ (28)
The transfer matrices $T_{ver}^{(\pm)}(x)$ are invariant under the $Z(3)$ spin rotation and their Hamiltonian limits, apart from a constant bulk term, correspond to the quantum spin chains $H^{(\pm)}$. Thus, we have the following commutation relations,
$T_{ver}^{(\pm)}(x),{\cal{O}}_{Z(3)}]=0,~~[T_{ver}^{(\pm)}(x),H^{(\pm)}]=0$ (29)
such that the spectra of $T_{ver}^{(+)}(x)$ and $T_{ver}^{(-)}(x)$ are related by the following equivalences,
$\mathrm{Spec}\big[T_{ver}^{(+)}\big]_{Q=0}=\mathrm{Spec}\big[T_{ver}^{(-)}\big]_{Q=0},~~\mathrm{Spec}\big[T_{ver}^{(+)}\big]_{Q=1}=\mathrm{Spec}\big[T_{ver}^{(-)}\big]_{Q=2},~~\mathrm{Spec}\big[T_{ver}^{(+)}\big]_{Q=2}=\mathrm{Spec}\big[T_{ver}^{(-)}\big]_{Q=1}.$ (30)
The transfer matrix $T_{ver}^{(c)}(x)$ keeps only the invariant concerning the $Z(2)$ subgroup associated with the charge conjugation. The respective Hamiltonian limit gives rise to the quantum spin chain $H^{(c)}$, namely
$T_{ver}^{(c)}(x),{\cal{O}}_{Z(2)}]=0,~~[T_{ver}^{(c)}(x),H^{(c)}]=0$ (31)
In the next sections, we shall see that the construction of these transfer matrices is essential to set up the Bethe equations for the corresponding quantum spin chains. From the spectral relations (30) it is enough to consider the problem for $T_{ver}^{(+)}(x)$ and $T_{ver}^{(c)}(x)$.
3 The Bethe equation for $T_{ver}^{(+)}(x)$
Here, we follow the strategy devised by Baxter to solve the eigenvalue problem associated with the diagonal-to-diagonal transfer matrix of the Ising model [1]. We recall that this kind of approach has also been used to uncover the Bethe equation for the critical $Z(n)$ spin model with periodic boundary conditions [11, 22]. The first step is to build a suitable proposal for the form of the eigenvalues of the transfer matrix $T_{ver}^{(+)}$. To this end, we observe any matrix elements of such operator have the following structure,
$\left[T_{ver}^{(+)}\right]_{a_{1},\dots,a_{L}}^{b_{1},\dots,b_{L}}=\Big[\frac{1}{g(x)g_{1}(x)}\Big]^{L}H(x)$ (32)
@@ 3 The Bethe equation for $T_{ver}^{(+)}(x)$
$T_{ver}^{(+)}(x)\left|\phi\right\rangle=\Lambda^{(+)}(x)\left|\phi\right\rangle$ (35)
@@
Due to the independence of the eigenstates on $x$ a given eigenvalue $\Lambda^{(+)}(x)$ can be regarded as a linear combination of the elements of the transfer matrix $T_{ver}^{(+)}(x)$ with constant coefficients. As a consequence, we can conclude that $\Lambda^{(+)}(x)$ should have the same form as given by Eqs.(32,34). To obtain further insights on the structure of $\Lambda^{(+)}(x)$ we have studied the eigenvalue problem (35) for small finite sizes. This analysis reveals to us that the the form of eigenvalues depends on the $Z(3)$ sectors besides not given solely in terms of the product of sines. Investigating the asymptotic behavior of $\Lambda^{(+)}(x)$ for $x\rightarrow\pm i\infty$ leads to propose the following form,
@@
The next step is to determine the set of nonlinear equations satisfied by the variables $\xi_{k}$ on a finite interval. This task can be done by exploring certain matrix identity among the products of transfer matrices with suitable distinct spectral parameters. Here we have been inspired by the matrix structure derived for the diagonal-to-diagonal transfer matrix of the three-state Chiral Potts model [12, 13]. We have found that the transfer matrix $T_{ver}^{(+)}(x)$ satisfies the following identity,
$T_{ver}^{(+)}(x-\frac{\pi}{3})T_{ver}^{(+)}(x-\frac{\pi}{6})T_{ver}^{(+)}(x)$ $=$ $T_{ver}^{(+)}(0)\bigg\{\big[f_{1}(x)\big]^{L}T_{ver}^{(+)}(x-\frac{\pi}{3})+\big[f_{2}(x)\big]^{L}T_{ver}^{(+)}(x)$ (38)
$+$ $\big[f_{3}(x)\big]^{L}T_{ver}^{(+)}(x+\frac{\pi}{3})\bigg\}$
@@
The way we have established the functional relation (38,39) is as follows. We first solve analytically some suitable entries of this matrix relation for $L=2$ and $L=3$ allowing the determination of the coefficients $f_{1}(x)$, $f_{2}(x)$ and $f_{3}(x)$. With the help of symbolic algebra packages we have verified that all the entries of the functional relation are indeed satisfied up to $L=5$. In this task we find convenient to transform the trigonometric functions in terms of ratio of polynomials in the variable $\exp(2ix)$. Based on this checking we conjecture that Eqs.(38,39) should be valid for arbitrary $L$ but a systematic proof of that has eluded us so far.
@@
At this point, we note that Eqs.(41,42) leads us to the expressions of the Bethe equation (11) and the energy eigenvalue (12) presented in the introduction by choosing,
@@ 3.1 The Bethe roots for $L=2$
We now describe the set of roots solving the Bethe equation (11) associated with the energy eigenvalues of the Hamiltonian $H^{(+)}$ for $L=2$. We obtain some insight on the nature of the roots, exploring the fact that the eigenvectors of $T_{ver}^{(+)}(x)$ are independent of the spectral parameter. By choosing a suitable value for $x$ we can compute numerically the eigenvectors making it possible the determination of the corresponding eigenvalues as polynomial functions on the variable $\exp(2ix)$. The zeros of such polynomials work as starting values for the numerical solution of the Bethe equation (11) providing us the Bethe roots with high precision.
@@
In table (1) we have summarized our results for $L=2$ considering the total number of nine eigenstates. For each energy of the spectrum of $H^{(+)}$ we exhibit the sector $Q$, the spin $s_{p}$ of the state and the corresponding roots $\{\lambda_{j}\}$ solving the Bethe equation (11). The energy values are obtained with the help of Eqs.(12,13) and we have verified that they agree with those obtained from the exact diagonalization of the Hamiltonian $H^{(+)}$. We remark that these Bethe roots allows us to compute the corresponding transfer matrix eigenvalues with the the help of Eqs.(37,43) for arbitrary values of the spectral parameter $x$. As an example we have added in the last column of table (1) the eigenvalues $\Lambda^{(+)}(x)$ at the point $x=\frac{3}{5}$ reproducing the values attained from the diagonalization of the operator $T_{ver}^{(+)}(\frac{3}{5})$. From table (1) we note that if a given energy in sector $Q=1$ is parametrized by a set of roots $\{\lambda_{j}\}$ the same energy is obtained in sector $Q=2$ using the opposite set of roots $\{-\lambda_{j}\}$. From our solution, we see that this property is valid for arbitrary $L$ in accordance with the fact that the Hamiltonian spectra for sectors $Q=1,2$ are the same since besides the spectral relation (30) we have that the invariance under reflection maps $H^{(+)}$ to $H^{(-)}$. We note that such spectral degeneracy does not apply to the transfer matrix eigenvalues due to the fact that we only have the spectral equivalence (30) for generic values of the spectral parameter. In addition, from Eqs.(13,44) we see that the momenta of the states with roots $\{\lambda_{j}\}$ and $\{-\lambda_{j}\}$ are however opposite to each other.
$E^{(+)}$ $Q$ $s_{p}$ $\lambda_{j}$ $\Lambda^{(+)}(\frac{3}{5})$
-4.93624921 0 0 $\pm 0.53202156i$ $0.95403776$
-2.30940107 1 $\frac{1}{3}$ $0.06315064\pm 0.54254935i,~0.19822180$ $\begin{array}[c]{c}1.16822213\\ +0.04919793i\end{array}$
-2.30940107 2 $\frac{1}{3}$ $0.06315064\pm 0.54254935i$0.19822180-0.19822180 $\begin{array}[c]{c}1.16822213\\ -0.04919793i\end{array}$
-0.84529946 1 $\frac{2}{3}$ $0.16532518\pm 0.50188873i$0.10277717-0.10277717 $\begin{array}[c]{c}1.32149657\\ -0.07557454i\end{array}$
-0.84529946 2 $\frac{2}{3}$ $0.16532518\pm 0.50188873i,~0.10277717$ $\begin{array}[c]{c}1.32149657\\ +0.07557454i\end{array}$
1.15470053 0 1 $0,~\frac{i\pi}{2}$ $1.54404018$
3.15470053 1 $\frac{2}{3}$ 0.245380170.245380170.02638725-0.026387250.64959867-0.64959867 $\begin{array}[c]{c}1.79532310\\ +0.025796612i\end{array}$
3.15470053 2 $\frac{2}{3}$ 0.24538017-0.245380170.02638725,0.649598670.02638725,~0.64959867 $\begin{array}[c]{c}1.79532310\\ -0.025796612i\end{array}$
3.78154867 0 0 $\pm 0.12258177$ $1.88075838$
@@
Table 1: The eigenvalues ($E^{(+)}$) of the Hamiltonian $H^{(+)}$ (10) for $L=2$. For each energy, we indicate the sector ($Q$), the spin of the state $(s_{p})$ and the values of the Bethe roots ($\lambda_{j}$). We have also added the respective transfer matrix eigenvalue $\Lambda^{(+)}(x)$ at $x=\frac{3}{5}$.
@@
4 The Bethe equation for $T_{ver}^{(c)}(x)$
Here, we use the same procedure explained in the previous section to determine the parameterization of the spectrum of Hamiltonian $H^{(c)}$ in terms of a Bethe ansatz equation. We therefore avoid the repetition of technical details and concentrate our presentation to the main results. In particular, we find out that the structure of the eigenvalues of $T_{ver}^{(c)}(x)$ can be given as follows,